バーチャルリアリティ技術 2 単位 (選択)

Virtual Reality  仁木登 · 教授/ システム創生工学専攻光システム工学コース光情報システム工学講座

【授業目的】バーチャルリアリティの基本的な技術を習得することを目的とする． ; 【授業概要】バーチャルリアリティの技術と人間の関わりを解説し，これらを構 ; 成する映像技術や計算技術，また高難度作業の応用例について後述する．こ ; の授業は工業に関する科目である． ; 【授業形式】 [形態] ; 【キーワード】 [キーワード] ; 【先行科目】 [先行科目] ; 【関連科目】 [関連科目] ; 【履修要件】画像処理，パターン認識を履修していることが望ましいが，受講し ; ていない学生にも理解できるように配慮する． ; 【履修上の注意】 [注意] ; 【到達目標】 ; 1. バーチャルリアリティ技術に必要な基礎事項を理解する． ; 2. 映像技術，計算技術，高難度作業に応用されている各種バーチャルリアリ ; ティ技術について理解する． ; 【授業計画】 ; 1. バーチャルリアリティ技術概論 ; 2. バーチャルリアリティ技術に関する数学的基礎 (画像処理) ; 3. バーチャルリアリティ技術に関する数学的基礎 (パターン認識) ; 4. コンピュータグラフィックス ; 5. 3 次元座標変換 ; 6. 幾何形状モデリング法 ; 7. 陰線消去 · 陰面消去法 ; 8. レンダリング法 ; 9. 実時間コンピュータグラフィックス ; 10. 動画像処理 ; 11. 画像センシング ; 12. 人工現実感システム ; 13. 複合現実感システム ; 14. バーチャルリアリティ技術の医療分野への応用 ; 15. バーチャルリアリティ技術の工業分野への応用 ; 16. 最終試験【成績評価基準】バーチャルリアリティ技術に関する課題レポート 60%，最終試験

40% 【教科書】各トピックスに関する参考書を紹介する．【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216813 【対象学生】 [対象学生] 【連絡先】 [連絡先] 【備考】授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習をしたうえで授業をうけることが，授業の理解と単位取得のために必要である．

Virtual Reality 2 units (selection)

Noboru Niki · Professor / Optical Materials and Devices, Optical Systems Engineering, Systems Innovation Engineering

Target⟩ Introduction to the fundamental concepts of virtual reality. Outline⟩ Virtual Reality is a combination of human interface, graphics, sensor

technology, high performance computing, and networking. It allows the user to interact with an artificial environment created by computers. Using these technologies, one has the feeling of total immersion in a new environment. Here, these technologies and several VR simulations are described.

Style⟩ [形態] �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ It is desired to finish a course of digital signal processing, image

processing, and pattern recognition . Notice⟩ [注意]

Goal⟩

1. To understand fundamentals of virtual reality.

2. To understand various techniques for realizing virtual environments.

Schedule⟩

1. 1. Introduction

2. Mathematical Fundamentals (Image Processing)

3. Mathematical Fundamentals (Pattern Recognition)

4. Computer Graphics

5. Three-dimensional Coordinate System

6. Geometrical Modeling

7. Hidden Line and Surface Removal Methods

8. Advanced Rendering Topics

9. Real Time Computer Graphics

10. Image Processing of Motion Pictures

11. Image Sensing Technology

12. Virtual Reality Systems

13. Augmented Reality Systems

14. Advanced Applications of Virtual Reality 1

15. Advanced Applications of Virtual Reality 2

16. Final Examination Evaluation Criteria⟩ Project Report 60%, Final Examination 40% Textbook⟩ Reference books are introduced to each topic. Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216813 Student⟩ [対象学生] Contact⟩ [連絡先] Note⟩ Preparation (2hrs) and Review (2hrs) are required to take this lecture (2hrs).